OpenStudy (anonymous):

can you check my answer please solve using the multiplication principle: 8x=-56 the solution is x = -7 is this correct?

OpenStudy (anonymous):

yes it is correct: 8x=-56 divide both sides by 8 u would get: 8x/8=-56/8 x=-7 so good work my friend

OpenStudy (anonymous):

thanks kushashwa could you find the slope if it exists; x=-5

OpenStudy (anonymous):

assuming this is the equation of a line. you set it up y=-x-5 y=mx+b Compare the two equations, slope is m, or coefficient of x.

OpenStudy (anonymous):

chagunanas therefore the slope is considered undefined

OpenStudy (anonymous):

Why undefined? m or coefficient of x is -1.

OpenStudy (anonymous):

someone else gave the answer as undefined since you are given and x coordinate only it goes straight up and down if you were given y coordinated the slope would be zero

OpenStudy (anonymous):

You didn't give me a lot of info; that is why I said 'assuming that is equation of line.' So x=-5 is an x intercept. The method I was using works if you have the equation of the line.

OpenStudy (anonymous):

x=-5 usually refers to a straight line along -5. In that case, a straight line, the slope is always 0.

OpenStudy (anonymous):

No, no, I got that wrong, it is a vertical line. So may be that guy is right.

OpenStudy (anonymous):

see the problem states only: x=-5 find the slope if it exists

OpenStudy (anonymous):

Yeah, consider a line going straight up in the air, that's what the guy means when he says undefined.